# Quantized neural network for complex hologram generation

### v1

@@ 1 Introduction
A major challenge in realizing practical CGH applications is the substantial computational demand required to achieve photorealistic images. CGH requires numerical light propagation from a 3D scene to compute a hologram. This calculation is highly non-local, meaning that a single geometric primitive affects many hologram pixels. Therefore, the computational cost typically scales with the number of hologram pixels and the complexity of the 3D scene. Additionally, photorealistic holographic images require physically accurate light propagation models and time-consuming iterative optimization, posing a trade-off between computational cost and image quality. Considerable effort has been put into developing fast and high-quality CGH calculations, and recent progress is summarized in [8]. Traditional CGH algorithms have been accelerated by approximating a light propagation model [9, 10], using efficient 3D scene representations [11, 12], caching pre-computed results [13, 14], and leveraging hardware accelerators [15, 16, 17, 18, 19]. Recently, the integration of neural networks for CGH has gained attention for its potential to overcome the trade-off between computational cost and image quality [20, 21, 22, 23, 24]. One seminal study is tensor holography, which uses a convolutional neural network (CNN) to efficiently approximate point-based light propagation [25]. Another notable study is HoloNet, which uses a U-Net architecture and achieves high-fidelity CGH optimization in real time [26, 27]. Although neural networks have enabled fast CGH algorithms, they currently rely on high-performance graphics processing units (GPUs) for real-time processing, leading to bulky systems and high power consumption. Deploying them on computationally limited embedded systems, such as AR headsets and automotive head-up displays, requires more efficient models to reduce computational cost, memory footprint, and power consumption.
@@ 2.2 Model architecture and training
Our model includes three refinements over the original model used in tensor holography. First, we used Hardtanh rather than Tanh because it is computationally less expensive and can be removed by appropriately setting the clipping range. Second, we added a depthwise convolution with batch normalization before concatenating the output of the residual block and the skip-connected input. This helps align the histograms of the two concatenated activations. Finally, we used ReLU6 rather than ReLU to suppress the growth of the residual block outputs, thereby mitigating the accuracy degradation after quantization.
@@ 2.3 Neural network quantization
The clipping ranges of weights can be computed statically, but those of activations vary for each input and can be determined either dynamically or statically. Dynamic quantization calculates the range for each activation during inference and quantizes it on the fly. By contrast, static quantization calculates the range using calibration data before inference to pre-quantize the activations. While dynamic quantization can achieve higher accuracy than static quantization, it adds computational overhead for calculating the clipping range. Additionally, numerous model deployment tools for embedded platforms only support statically quantized models and do not accommodate dynamically quantized models.
@@
We quantized our FP32 model using post-training dynamic quantization (PTDQ), post-training static quantization (PTSQ), and static QAT. The quantization framework was PyTorch FX Graph Mode Quantization. The target backend was QNNPACK, and its default settings were used for calibration. For QAT, we further trained the pre-trained FP32 models for 20 epochs with a reduced learning rate of $1\times 10^{-6}$ and statically quantized the weights and activations.
@@ 3.1 Hologram quality
We evaluated the image quality of output holograms from the FP32 and INT8 models. The image quality was assessed using the peak signal-to-noise ratio (PSNR) and the structural similarity index measure (SSIM). These metrics were applied to both the amplitude and phase images of the complex hologram. Table 1 shows the average hologram quality for each model over the test dataset. In FP32 precision, both the original tensor holography model and our refined model had similar quality. However, the original INT8 model with PTSQ showed a significant decrease in hologram quality compared to FP32. Moreover, QAT did not improve this PTSQ model, suggesting the original model has architectural limitations that result in a significant drop in accuracy after static quantization. In contrast, all refined INT8 models achieved comparable hologram quality to that of FP32. Thus, our results demonstrate that the decrease in hologram quality caused by INT8 static quantization can be minimized through our architectural enhancements.
Table 1: Hologram quality of the original tensor holography model and our refined model in FP32 and INT8 with PTDQ, PTSQ, and QAT. PSNR and SSIM are averaged over the test dataset.
Model Precision Method PSNR (dB) SSIM
Amplitude Phase Amplitude Phase
Original FP32 - 29.2 28.3 0.907 0.903
Refined FP32 - 30.1 28.3 0.915 0.904
Original INT8 PTDQ 29.2 28.3 0.907 0.903
Refined INT8 PTDQ 30.1 28.3 0.915 0.904
Original INT8 PTSQ 7.78 24.3 0.205 0.555
Refined INT8 PTSQ 29.3 27.9 0.899 0.888
Original INT8 QAT 12.6 21.8 0.241 0.548
Refined INT8 QAT 29.2 28.2 0.904 0.897
@@ 3.2 Reconstructed images
For qualitative evaluation of hologram quality, we reconstructed images using the complex holograms output from the FP32 and INT8 models in simulation. Figure 2 shows the reconstructed images from the output complex holograms. The resolution of the input RGB-D data was 384 × 384 pixels. The angular spectrum method was used to compute the reconstructed images 6 mm away from the hologram plane. The results show that the original INT8 model with PTDQ accurately reconstructs the images, while those with PTSQ and QAT completely failed to do so. However, even after static quantization, our refined INT8 models achieved comparable image quality to the non-quantized FP32 model.
Figure 2: Reconstructed images from complex holograms computed by original tensor holography model and our refined model in FP32 and INT8 with PTDQ, PTSQ, and QAT. ©2008, Blender Foundation / www.bigbuckbunny.org.

### v2

@@ 1 Introduction
A major challenge in realizing practical CGH applications is the substantial computational demand required to achieve photorealistic images. CGH requires numerical light propagation from a 3D scene to compute a hologram. This calculation is highly non-local, meaning that a single geometric primitive affects many hologram pixels. Therefore, the computational cost typically scales with the number of hologram pixels and the complexity of the 3D scene. Additionally, photorealistic holographic images require physically accurate light propagation models and time-consuming iterative optimization, posing a trade-off between computational cost and image quality. Considerable effort has been put into developing fast and high-quality CGH calculations, and recent progress is summarized in [8]. Traditional CGH algorithms have been accelerated by approximating a light propagation model [9, 10], using efficient 3D scene representations [11, 12], caching pre-computed results [13, 14], and leveraging hardware accelerators [15, 16, 17, 18, 19]. Recently, the integration of neural networks for CGH has gained attention for its potential to overcome the trade-off between computational cost and image quality [20, 21, 22, 23, 24, 25, 26]. One seminal study is tensor holography, which uses a convolutional neural network (CNN) to efficiently approximate point-based light propagation [27]. Another notable study is HoloNet, which uses a U-Net architecture and achieves high-fidelity CGH optimization in real time [28, 29]. Although neural networks have enabled fast CGH algorithms, they currently rely on high-performance graphics processing units (GPUs) for real-time processing, leading to bulky systems and high power consumption. Deploying them on computationally limited embedded systems, such as AR headsets and automotive head-up displays, requires more efficient models to reduce computational cost, memory footprint, and power consumption.
@@ 2.2 Model architecture and training
Our model includes three refinements over the original model used in tensor holography. First, we used Hardtanh rather than Tanh because it is computationally less expensive and can be removed by appropriately setting the clipping range. Second, we added a depthwise convolution with batch normalization before concatenating the output of the residual block and the skip-connected input. This helps align the histograms of the two concatenated activations. Finally, we used ReLU6 rather than ReLU to suppress the growth of the residual block outputs, thereby mitigating the accuracy degradation after quantization. Figure 2 shows the histograms of two activations that are concatenated in both the original and refined models. This example clearly demonstrates that the original model introduces an imbalance in the histograms between the skip-connected input and the output of the final residual block. In contrast, our refined model addresses this issue by aligning the histograms, making them more consistent.
Figure 2: Histograms of two activations that are concatenated in both the original and refined models.
@@ 2.3 Neural network quantization
The clipping ranges of weights can be computed statically, but those of activations vary for each input and can be determined either dynamically or statically. Dynamic quantization calculates the range for each activation during inference and quantizes it on the fly. By contrast, static quantization calculates the range using a representative dataset (i.e., calibration dataset) before inference to pre-quantize the activations. While dynamic quantization can achieve higher accuracy than static quantization, it adds computational overhead for calculating the clipping range. Additionally, numerous model deployment tools for embedded platforms only support statically quantized models and do not accommodate dynamically quantized models.
@@
We quantized our FP32 model using PyTorch FX Graph Mode Quantization [34]. The quantization methods were post-training dynamic quantization (PTDQ), post-training static quantization (PTSQ), and static QAT, which are supported by the quantization framework. PTDQ and PTSQ are PTQ methods using dynamic and static quantization, respectively. The target backend was QNNPACK, and its default settings were used for calibration. The calibration dataset for PTSQ consisted of 100 samples from the training dataset. For QAT, we further trained the pre-trained FP32 models for 20 epochs with a reduced learning rate of $1\times 10^{-6}$ and statically quantized the weights and activations.
@@ 3.1 Hologram quality
We evaluated the image quality of output holograms from the FP32 and INT8 models. The image quality was assessed using the peak signal-to-noise ratio (PSNR) and the structural similarity index measure (SSIM). These metrics were applied to both the amplitude and phase images of the complex hologram. Figure 3 shows the average hologram quality for each model over the test dataset. In FP32 precision, both the original tensor holography model and our refined model had similar quality. In INT8 precision, both the original and refined PTDQ models produced holograms nearly identical to those of the FP32 models. However, the original PTSQ model showed a significant decrease in hologram quality compared to FP32. Moreover, QAT did not improve this PTSQ model, suggesting the original model has architectural limitations that result in a significant drop in accuracy after static quantization. In contrast, all refined INT8 models achieved comparable hologram quality to that of FP32. The improvement with QAT was minimal since the refined PTSQ model already exhibited sufficient hologram quality. As shown in Section 2.2, the significant accuracy drop in the original model is mainly caused by an imbalance in the histograms between the skip-connected input and the output of the final residual block. Our refined model addresses this issue by using ReLU6 to suppress the growth of the residual block outputs and by aligning the histograms through the convolution layers prior to concatenation. These results demonstrate that the decrease in hologram quality caused by INT8 static quantization can be minimized through our architectural enhancements.
Figure 3: Hologram quality of the original tensor holography model and our refined model in FP32 and INT8 with PTDQ, PTSQ, and QAT. PSNR and SSIM are averaged over the test dataset.
@@ 3.2 Reconstructed images
We reconstructed images using the complex holograms output from the FP32 and INT8 models in simulation. Figure 4 shows the reconstructed images from the output complex holograms. The resolution of the input RGB-D data was 384 × 384 pixels. The angular spectrum method was used to compute the reconstructed images 2 mm away from the hologram plane. We computed the PSNR and SSIM between the reconstructed images from the INT8 and FP32 models to evaluate the impact of quantization. The results show that the original INT8 model with PTDQ accurately reconstructs the images, while those with PTSQ and QAT completely failed to do so. However, even after static quantization, our refined INT8 models achieved comparable image quality to the non-quantized FP32 model.
Figure 4: Reconstructed images from complex holograms computed by original tensor holography model and our refined model in FP32 and INT8 with PTDQ, PTSQ, and QAT. The PSNR and SSIM are computed between the reconstructed images from the INT8 and FP32 models. ©2008, Blender Foundation / www.bigbuckbunny.org.